OpenStudy (anonymous):

Which of the following equations has solutions of 3 ± sqrt 5? a. (x - 5)^2 = 3 b. (x - 3)^2 = 5 c. (x + 3)2 = 5 d. (x + 5)^2 = 3

OpenStudy (anonymous):

oops, c. (x + 3)^2 = 5

OpenStudy (jdoe0001):

let's test the 1st one \(\bf (x - 5)^2 = 3\\ \textit{let's get square root on both sides}\\ \pm\sqrt{ (x - 5)^2 } = \pm\sqrt{ 3 } \implies x - 5 = \sqrt{ 3 }\\ \textit{add 5 to both sides}\\ x \cancel{- 5 + 5} = 5\pm\sqrt{ 3 } \implies x = 5\pm\sqrt{ 3 }\) so, that's not it, then try the 2nd one